# The influence of substantial intragranular orientation gradients on the micromechanical response of heavily-worked material

### v1

@@ Abstract
In this study, we examine the role of relatively large amounts of intragranular lattice misorientation – present after many thermomechanical processes – through three-dimensional (3D) micromechanical simulations of forged Al-7085 with a modified T7452 temper. We utilize near-field high-energy X-ray diffraction microscopy (HEDM) to measure 3D spatial orientation fields, and employ a novel reconstruction method that utilizes grain orientation envelopes measured using far-field HEDM to enable reconstruction of grains with intragranular orientation spreads greater than . We construct two primary virtual polycrystalline specimens for use in crystal plasticity simulations to assess the effect of appreciable intragranular misorientation on the predicted deformation response: the first a faithful representation of the HEDM reconstruction, the second a microstructure with no intragranular misorientation (i.e., grain-averaged orientations). We find significant differences in the predicted deformation mechanism activation, distribution of stress, and distribution of plastic strain between simulations containing intragranular misorientation and those with homogenized orientations. The influence of elastic anisotropy is discussed, along with the effects of intragranular misorientation on fatigue life through the calculation and analysis of fatigue indicator parameters.
@@ 1 Introduction
In addition to grain reorientation and macroscopic texture evolution that occurs during thermomechanical processing, there is generally an accompanying development of crystallographic misorientation within individual grains [1, 2]. Similarly, intragranular misorientation develops during many solidification processes [3, 4], including during most additive manufacturing processes [5, 6]. The magnitudes of intragranular misorientation present have the potential to be significant, on the order of 10’s of degrees. As local orientation is highly influential on the micromechanical response, heterogeneous intragranular orientation fields have the potential to lead to large variations in mechanical state and active plastic deformation mechanisms. However, the challenge with modeling—and ultimately quantifying—these effects lies with the ability to instantiate a model with realistic intragranular orientation fields (i.e., including intragranular misorientation). This misorientation present within grains results from underlying three-dimensional (3D) geometrically necessary dislocation (or Nye) tensor fields, which are described by a very large parameter space that is not readily evaluated. A more tractable approach is to instantiate simulations with experimentally measured orientation fields which, while not capturing the full parameter space, do reflect realistic microstructural conditions. However, means to collect 3D orientation fields, particularly in heavily deformed materials, are limited. In this study, we present a new methodology for reconstructing 3D orientation fields with large intragranular orientation gradients using high-energy X-ray diffraction microscopy (HEDM) and apply this methodology to forged Al-7085. The data is used to instantiate crystal plasticity finite element method (CPFEM) simulations and explore what role orientation fields with appreciable intragranular misorientation play in (micro)mechanical response.
@@
In this paper, we utilize a 3D microstructure reconstructed using data from nf-HEDM and ff-HEDM to explore how large orientation gradients that arise during metal processing influence micromechanical response. First, we describe the experimental and novel reconstruction methodology used to reconstruct 3D orientation fields with the presence of appreciable intragranular orientation gradients in Al-7085. Next, we describe the instantiation of the virtual polycrystals for use in crystal plasticity simulations, in particular the instantiation of the faithful representation of the experimentally-deduced polycrystal, as well as the grain-averaged homogenized polycrystal. Next, results comparing the predicted micromechanical response of these two samples are presented. Lastly, we discuss the relationships between intragranular orientation gradients and plastic deformation mechanism activation, the influence of elastic anisotropy, and finally the fatigue response.
@@ 2.1 Material and Measurement Description
The incoming X-ray beam had an energy of (Pr Kα-edge) and a width of to fully illuminate the sample cross section. The beam height was , and we probed 4 volumes with a spacing of (the vertically oversized beam facilitates stitching and allows data at the beam edges to be discarded). In total, we probed a volume spanning along L, along LT, and along ST. For both nf-HEDM and ff-HEDM, we collected diffraction data across a full in increments (1440 images total) as we continuously rotated the specimen about the ST direction. Prior to the HEDM measurements, we performed an XCT scan using a wide by tall beam with X-ray absorption radiographs made at increments over . These XCT measurements facilitate the establishment of the physical extent of the specimen to speed data reconstruction.
@@ 2.2 Data Processing and Microstructure Reconstructions
Following this methodology, we perform indexing on the far-field diffraction data using the trial orientation grid. We employ a completeness threshold of 0.79, selected from direct analysis of the full distribution of completeness values (i.e., we selected a threshold necessary to index the majority of the material volume without incorporating spurious orientation solutions). The lattice parameter () used for predicting diffracted intensity positions was . After indexing all four volumes probed during measurement, we concatenate the lists of individual orientations from each volume into a global list for orientation testing. In total, we found 181,300 orientations in the full diffraction volume, and present a plot of these orientations in the cubic symmetry fundamental region of Rodrigues space (using the Rodrigues orientation parameterization, , and a crystal-to-sample convention) in Figure 2a. We observe distinct, closed “envelopes” of orientations (in other words: contiguous point-clouds), which we will herein refer to as grain orientation envelopes (GOEs). We note that these GOEs generally correspond to grains within the diffraction volume (i.e., the orientations within GOEs are contiguous in real-space), indicating that they do indeed belong to single grains. To isolate these different envelopes, we employ the DBSCAN clustering algorithm [27] with parameters of (the clustering distance) and 32 orientations for the minimum cluster size, with parameters chosen to isolate envelopes that were visually apparent by inspection. After clustering, we manually identified envelopes which are discontinuous in orientation space due to crystal symmetry (i.e., exit one side and appear on the other, see [28]) and combine them. The GOEs which reside on symmetric equivalencies of the fundamental region boundaries tend to have similar point cloud shapes on the boundaries, lending confidence that they indeed belong to the same GOE and are separated only due to crystal symmetry. After this procedure, we found 17 distinct GOEs, as shown in Figure 2b (colored by GOE). We note the presence of significant variation in GOE size in orientation space, which is particularly noteworthy as several of these GOEs indicate granular orientation spreads on the order of to .
@@ 3.1 Crystal Plasticity Model and Material Parameters
We performed micromechanical simulations using the FEPX CPFEM package. Here we provide only a truncated description of the model employed in FEPX below. A complete implementation is detailed in [29, 30, 31, 32]. The FEPX framework considers ductile, isothermal, quasi-static material behavior, and employs a non-linear finite element solver to approximate the solution to the equilibrium via an implicit formulation to force the reduction of a global weighted residual (thus ensuring the approximate enforcement of stress equilibrium). In this formulation, the lattice orientation and slip system strength are stored as evolving state variables at the quadrature points of each element.
@@
where is the stress projected onto slip system (i.e., the resolved shear stress), is the current slip system strength (or hardness), and is the power law exponent. The resolved shear stress on each slip system is calculated as
where is the deviatoric portion of the stress. We allow the current slip system strength to evolve as a function of local plasticity. We employ an isotropic (i.e., the single crystal yield surface may dilate, but retains its shape in deviatoric stress space) saturation-style hardening model:
@@
We will often be analyzing the total local plastic strain which we define as
@@ 3.2 Virtual Microstructure Instantiation and Loading
Overall, we choose to instantiate two separate virtual microstructures for use in the CPFE simulations in an effort to determine the effects of incorporating appreciable intragranular orientation spreads: the first, a faithful representation of the spatial orientation field, and the second, a structure with grain-averaged orientations applied homogeneously within each grain. In both structures, we utilize the nf-HEDM reconstruction in Figure 3a to instantiate the spatial location of the grains in the sample. In this way, we faithfully represent the geometric morphology of the granular structure in the virtual samples. To achieve this, we construct a regular grid of voxels and perform a direct mapping of the measured orientations of the experimental microstructure to that of the virtual grid. This virtual grid may have coarser or finer resolution than that of the experimental grid, depending on desired model resolution. In this study, due to computational constraints, we will utilize a virtual polycrystal with a coarser voxel grid than the experimental grid. In this case, we assign an orientation to a point in a coarse grid of the virtual polycrystal based on the dominant experimental orientation in its direct vicinity (i.e., the most prevalent orientation within the element bounds).
@@
In the first of the two microstructures (i.e., a faithful representation of the spatial orientation field), we apply local orientations utilizing a similar scheme above employed for the assignment of grain ID. Here, a point in a coarse grid of the virtual polycrystal is assigned an orientation based on the average orientation of the experimental voxels with common grain ID in the local vicinity using orientation averaging functionality in the HEXRD package that accounts for crystal symmetry. In this way, we ensure that the orientation field of the virtual polycrystal matches that measured experimentally. In the second of the two microstructures, we assign a single orientation to each voxel with a common grain ID, where we calculate this orientation to be equal to the average of all orientations in the experimental dataset with the same grain ID (i.e., we average over the GOEs shown in Figure 2a). This method is consistent with the approaches of the overwhelming majority of crystal plasticity studies, and obscures the gradient of orientations present within grains. From this point onwards, we will refer to the first of the two microstructures as the “gradient” structure, and the second of the two as the “homogenized” structure. In both the gradient and homogenized microstructures, each voxel in the grid is split into 6, 10-node tetrahedral elements, with each element assigned the same grain ID and orientation as its parent voxel.
@@
Following the general procedure above, we begin by first removing a thin layer of material on the experimental data set associated with the rough sample surface. We reduced the sample size from 1 0.5 to 0.9 0.45 due to this truncation process (i.e., the reduced volume is left comprised of 180 90 180 voxels). We further perform de-noising of the nf-HEDM orientation field by finding isolated volumes of orientation with sizes less than 64 voxels and then setting these orientations to match the neighboring region (compare: Figure 2a and b, and note the removal of small orientation features as a consequence of this de-noising). We map the nf-HEDM microstructure to the virtual polycrystal grid with mesh coarsening of 2—i.e., we concatenate 8 experimental voxels (2 2 2) into a single voxel in the virtual polycrystal grid, resulting in a virtual polycrystal with a voxel grid of 90 45 90 voxels and thus reducing the mesh size 8 compared to a direct 1-to-1 map. Considering the 6 tetrahedral elements per voxel, the overall mesh contains approximately 2.2 million elements. We present the gradient and homogenized virtual microstructures in Figure 4a and Figure 4b respectively.
@@
We applied loading to the two virtual microstructures along their ST () directions to explore the variation of micromechanical response due to the differences in orientation present in the material. We load through the yield point and into fully-developed plastic flow to macroscopic engineering strains of 0.2. We employ boundary conditions via velocities applied at select nodes of the finite element mesh. We apply minimal boundary conditions to the sample, where the domain is constrained to apply axial deformation yet disallow rigid body rotation or motion, by applying axial velocities on two surfaces, and fixing two nodes in orthogonal directions. This, overall, allows the polycrystal to deform with minimal undue influence from the boundary conditions which may over-constrain portions of the domain and lead to stress concentrations and over-predictions. On the loading surface (normal along ), we apply velocities on the control surface to enforce a constant engineering strain rate of .
@@ 4.1 Macroscopic Behavior
In this section, we contrast the mechanical response of the gradient microstructure versus the homogenized microstructure. We present the macroscopic stress-strain (-) response up to engineering strains of 0.2 (20%) for both microstructures in Figure 5. We will provide more detailed analysis at 4 points along the loading history, marked in Figure 5, specifically at macroscopic engineering strains of 0.003 (elastic regime), 0.011 (yield), 0.08 (post-yield), and 0.2 (fully-developed flow). Again, we note that the same elasticity and plasticity parameters (first row of Table 1, and Table 2, respectively) are utilized when simulating both the gradient and homogenized microstructures—i.e., the simulations differ only in microtexture / localized distributions of orientations. While the yield stresses of the two microstructures are nearly identical (see Table 3), we observe that the flow stresses of the two virtual microstructures begins to deviate significantly from one another shortly after macroscopic yield and the differences continue to grow as plasticity develops (e.g., at , the flow stress of the gradient microstructure is over higher than that of the homogenized microstructure). The macroscopic behavior of the gradient microstructure exhibits an appreciably higher macroscopic hardening rate, and does not reach a point of macroscopic saturation (i.e., a peak in the stress-strain curve). Conversely, the macroscopic behavior of the homogenized microstructure exhibits a lower macroscopic hardening rate, and is approaching the peak of its stress-strain behavior.
@@ 4.2 Element-Scale Stress Behavior
We can further view these differences between the predicted von Mises stresses on an element-by-element basis between the two polycrystals. We present this data in Figure 7, which depicts two-dimensional histograms comparing the element-by-element von Mises stresses between the gradient and homogenized virtual microstructures at the same four macroscopic states of interest. For interpreting the figures, we note that the diagonal line indicates perfect correlation (“1-to-1”)—in other words, the histogram bins which fall on this line indicate the counts of elements which have the same von Mises stress in both the gradient simulation and the homogenized simulation. Histogram bins above the line indicate counts of elements which had a higher von Mises stress in the gradient sample, while those below the line indicate the opposite. White regions in the plots are bins that have no counts of elements which exhibit the given stress pairing between the gradient and homogenized microstructures. We observe, qualitatively, that as plasticity begins and evolves, the distribution of element counts tends to skew toward a positive deviation from the line of perfect correlation, indicating that more elements exhibit higher von Mises stresses in the gradient virtual microstructure compared to the homogenized microstructure.

### v2

@@ Abstract
In this study, we employ high-energy X-ray characterization to examine the role of relatively large amounts of intragranular lattice misorientation – present after many thermomechanical processes – on the micromechanical response of Al-7085 with a modified T7452 temper. We utilize near-field high energy X-ray diffraction microscopy (HEDM) to measure three-dimensional (3D) spatial orientation fields, facilitated by a novel method that utilizes grain orientation envelopes measured using far-field HEDM to enable reconstruction of grains with intragranular orientation spreads greater than . We then assess the consequences of consideration of intragranular orientation fields on the predicted deformation response of the sample through 3D micromechanical simulations of the forged Al-7085. We construct two virtual polycrystalline specimens for use in simulations: the first a faithful representation of the HEDM reconstruction, the second a microstructure with no intragranular misorientation (i.e., grain-averaged orientations). We find significant differences in the predicted deformation mechanism activation, distribution of stress, and distribution of plastic strain between simulations containing intragranular misorientation and those with grain-averaged orientations, indicating the necessity for consideration of intragranular orientation fields for accurate predictions. Further, the influence of elastic anisotropy is discussed, along with the effects of intragranular misorientation on fatigue life through the calculation and analysis of fatigue indicator parameters.
@@ 1 Introduction
In addition to grain reorientation and macroscopic texture evolution that occurs during thermomechanical processing, there is generally an accompanying development of crystallographic misorientation within individual grains [1, 2]. Similarly, intragranular misorientation develops during many solidification processes [3, 4], including during most additive manufacturing [5, 6]. The magnitudes of intragranular misorientation present have the potential to be significant, on the order of 10’s of degrees. As local orientation is highly influential on the micromechanical response, heterogeneous intragranular orientation fields have the potential to lead to large variations in mechanical state and active plastic deformation mechanisms. However, the challenge with micromechanical modeling—and ultimately quantifying—these effects lies with the ability to instantiate a model with realistic intragranular orientation fields (i.e., including intragranular misorientation). This misorientation present within grains results from underlying three-dimensional (3D) geometrically necessary dislocation (or Nye) tensor fields, which are described by a very large parameter space that is not readily evaluated. A more tractable approach is to instantiate simulations with experimentally measured orientation fields which, while not capturing the full parameter space, do reflect realistic microstructural conditions. However, means to collect 3D orientation fields, particularly in heavily deformed materials, are limited and often destructive [7], the latter eliminating possibility of further in situ micromechanical experiments on the measured sample. To address this issue, we present a novel methodology to non-destructively reconstruct 3D orientation fields with large intragranular orientation gradients using high-energy X-ray diffraction microscopy (HEDM), and demonstrate this methodology on a forged Al-7085 sample. The data is used to instantiate crystal plasticity finite element method (CPFEM) simulations and explore what role orientation fields with appreciable intragranular misorientation play in (micro)mechanical response.
@@
We present a novel methodology to reconstruct the spatial fields of orientations in 3D samples with large orientation gradients using a combined nf-HEDM and ff-HEDM approach, and further explore how large orientation gradients that arise during metal processing influence micromechanical response via a suite of crystal plasticity simulations. First, we describe the experimental and novel reconstruction methodology used to reconstruct 3D orientation fields with the presence of appreciable intragranular orientation gradients in Al-7085. Next, we describe the instantiation of the virtual polycrystals for use in crystal plasticity simulations, in particular the instantiation of the faithful representation of the experimentally-deduced polycrystal, as well as the grain-averaged homogenized polycrystal. Next, results comparing the predicted micromechanical response of these two samples are presented. Lastly, we discuss the relationships between intragranular orientation gradients and plastic deformation mechanism activation, the influence of elastic anisotropy, and finally the fatigue response.
@@ 2.1 Material and Measurement Description
We utilized an incoming X-ray beam with an energy of (Pr Kα-edge) and a width of to fully illuminate the sample cross section. We utilized a beam height of , and we probed 4 volumes with a spacing of (the vertically oversized beam facilitates stitching and allows data at the beam edges to be discarded). In total, we probed a volume spanning along L, along LT, and along ST. For both nf-HEDM and ff-HEDM, we collected diffraction data across a full in increments (1440 images total) as we continuously rotated the specimen about the ST direction. Prior to the HEDM measurements, we performed an XCT scan using a wide by tall beam with X-ray absorption radiographs made at increments over . These XCT measurements facilitate the establishment of the physical extent of the specimen to speed data reconstruction.
@@ 2.2 Data Processing and Microstructure Reconstructions
Following this methodology, we perform indexing on the far-field diffraction data using the trial orientation grid. We employ a completeness threshold of 0.79, selected from direct analysis of the full distribution of completeness values (i.e., we selected a threshold necessary to index the majority of the material volume without incorporating spurious orientation solutions). We utilize a lattice parameter () of for predicting diffracted intensity positions. After indexing all four volumes probed during measurement, we concatenate the lists of individual orientations from each volume into a global list for orientation testing. In total, we find 181,300 orientations in the full diffraction volume, and present a plot of these orientations in the cubic symmetry fundamental region of Rodrigues space (using the Rodrigues orientation parameterization, , and a crystal-to-sample convention) in Figure 2a. We observe distinct, closed “envelopes” of orientations (in other words: contiguous point-clouds), which we will herein refer to as grain orientation envelopes (GOEs). We note that these GOEs generally correspond to grains within the diffraction volume (i.e., the orientations within GOEs are contiguous in real-space), indicating that they do indeed belong to single grains. To isolate these different envelopes, we employ the DBSCAN clustering algorithm [27] with parameters of (the clustering distance) and 32 orientations for the minimum cluster size, with parameters chosen to isolate envelopes that were visually apparent by inspection. After clustering, we manually identified envelopes which are discontinuous in orientation space due to crystal symmetry (i.e., exit one side and appear on the other, see [28]) and combine them. The GOEs which reside on symmetric equivalencies of the fundamental region boundaries tend to have similar point cloud shapes on the boundaries, lending confidence that they indeed belong to the same GOE and are separated only due to crystal symmetry. We find 17 distinct GOEs, as shown in Figure 2b (colored by GOE), and note the presence of significant variation in GOE size in orientation space, which is particularly noteworthy as several of these GOEs indicate granular orientation spreads on the order of to .
@@ 3.1 Crystal Plasticity Model and Material Parameters
We perform micromechanical simulations using the FEPX CPFEM package. Here we provide only a truncated description of the model employed in FEPX below. A complete implementation is detailed in [29, 30, 31, 32]. The FEPX framework considers ductile, isothermal, quasi-static material behavior, and employs a non-linear finite element solver to approximate the solution to the equilibrium via an implicit formulation to force the reduction of a global weighted residual (thus ensuring the approximate enforcement of stress equilibrium). In this formulation, the lattice orientation and slip system strength are stored as evolving state variables at the quadrature points of each element.
@@
where is the stress projected onto slip system (i.e., the resolved shear stress), is the current slip system strength (or hardness), and is the power law exponent. We calculate the resovled shear stress on each slip system as:
where is the deviatoric portion of the stress. We allow the current slip system strength, , to evolve as a function of local plasticity. We employ an isotropic (i.e., the single crystal yield surface may dilate, but retains its shape in deviatoric stress space) saturation-style hardening model:
@@
To aid in analysis of the predicted deformation fields, we will analyze the total local plastic strain, , which we define as:
@@ 3.2 Virtual Microstructure Instantiation and Loading
Overall, we choose to instantiate two separate virtual microstructures for use in the CPFE simulations in an effort to determine the effects of incorporating appreciable intragranular orientation spreads: the first, a faithful representation of the spatial orientation field, and the second, a structure with grain-averaged orientations applied homogeneously within each grain. In both structures, we utilize the nf-HEDM reconstruction in Figure 3a to instantiate the spatial location of the grains in the sample. In this way, we faithfully represent the geometric morphology of the granular structure in the virtual samples. To achieve this, we construct a regular grid of voxels and perform a direct mapping of the measured orientations of the experimental microstructure to that of the virtual grid. This virtual grid may have coarser or finer resolution than that of the experimental grid, depending on desired model resolution. In this study, due to computational constraints, we utilize a virtual polycrystal with a coarser voxel grid than the experimental grid. In this case, we assign an orientation to a point in a coarse grid of the virtual polycrystal based on the dominant experimental orientation in its direct vicinity (i.e., the most prevalent orientation within the element bounds).
@@
In the first of the two microstructures (i.e., a faithful representation of the spatial orientation field), we apply local orientations utilizing a similar scheme above employed for the assignment of grain ID. Here, a point in a coarse grid of the virtual polycrystal is assigned an orientation based on the average orientation of the experimental voxels with common grain ID in the local vicinity using orientation averaging functionality in the HEXRD package that accounts for crystal symmetry. In this way, we ensure that the orientation field of the virtual polycrystal matches that measured experimentally. In the second of the two microstructures, we assign a single orientation to each voxel with a common grain ID, where we calculate this orientation to be equal to the average of all orientations in the experimental dataset with the same grain ID (i.e., we average over the GOEs shown in Figure 2a). This method is consistent with the approaches of the overwhelming majority of crystal plasticity studies, and obscures the gradient of orientations present within grains. From this point onwards, we refer to the first of the two microstructures as the “gradient” structure, and the second of the two as the “homogenized” structure. In both the gradient and homogenized microstructures, each voxel in the grid is split into 6, 10-node tetrahedral elements, with each element assigned the same grain ID and orientation as its parent voxel.
@@
Following the general procedure above, we begin by first removing a thin layer of material on the experimental data set associated with the rough sample surface. We reduce the sample size from 1 0.5 to 0.9 0.45 due to this truncation process (i.e., the reduced volume is left comprised of 180 90 180 voxels). We further perform de-noising of the nf-HEDM orientation field by finding isolated volumes of orientation with sizes less than 64 voxels and then setting these orientations to match the neighboring region (compare: Figure 2a and b, and note the removal of small orientation features as a consequence of this de-noising). We map the nf-HEDM microstructure to the virtual polycrystal grid with mesh coarsening of 2—i.e., we concatenate 8 experimental voxels (2 2 2) into a single voxel in the virtual polycrystal grid, resulting in a virtual polycrystal with a voxel grid of 90 45 90 voxels and thus reducing the mesh size 8 compared to a direct 1-to-1 map. Considering the 6 tetrahedral elements per voxel, the overall mesh contains approximately 2.2 million elements. We present the gradient and homogenized virtual microstructures in Figure 4a and Figure 4b respectively. The grain reference orientation deviation (GROD) between the two microstructures calculated on an element-by-element basis is shown in Figure 4c.
@@
We apply loading to the two virtual microstructures along their ST () directions to explore the variation of micromechanical response due to the differences in orientation present in the material. We load through the yield point and into fully-developed plastic flow to macroscopic engineering strains of 0.2. We employ boundary conditions via velocities applied at select nodes of the finite element mesh. We apply minimal boundary conditions to the sample, where the domain is constrained to apply axial deformation yet disallow rigid body rotation or motion, by applying axial velocities on two surfaces, and fixing two nodes in orthogonal directions. This, overall, allows the polycrystal to deform with minimal undue influence from the boundary conditions which may over-constrain portions of the domain and lead to stress concentrations and over-predictions. On the loading surface (normal along ), we apply velocities on the control surface to enforce a constant engineering strain rate of .
@@ 4.1 Macroscopic Behavior
In this section, we contrast the mechanical response of the gradient microstructure versus the homogenized microstructure. We present the macroscopic stress-strain (-) response up to engineering strains of 0.2 (20%) for both microstructures in Figure 5. We provide more detailed analysis at 4 points along the loading history, marked in Figure 5, specifically at macroscopic engineering strains of 0.003 (elastic regime), 0.011 (yield), 0.08 (post-yield), and 0.2 (fully-developed flow). Again, we note that we utilize the same elasticity and plasticity parameters (first row of Table 1, and Table 2, respectively) when simulating both the gradient and homogenized microstructures—i.e., the simulations differ only in microtexture / spatial distributions of orientations. While the yield stresses of the two microstructures are nearly identical (see Table 3), we observe that the flow stresses of the two virtual microstructures begin to deviate significantly from one another shortly after macroscopic yield and the differences continue to grow as plasticity develops (e.g., at , the flow stress of the gradient microstructure is over higher than that of the homogenized microstructure). The macroscopic behavior of the gradient microstructure exhibits an appreciably higher macroscopic hardening rate, and does not reach a point of macroscopic saturation (i.e., a peak in the stress-strain curve). Conversely, the macroscopic behavior of the homogenized microstructure exhibits a lower macroscopic hardening rate, and is approaching the peak of its stress-strain behavior. These differences in macroscopic strength and hardening rate are a result of the differences in local microtexture between the two virtual specimens rather than a size effect.
@@ 4.2 Element-Scale Stress Behavior
We can further view these differences between the predicted von Mises stresses on an element-by-element basis between the two polycrystals. We present this data in Figure 7, which depicts two-dimensional histograms comparing the element-by-element von Mises stresses between the gradient and homogenized virtual microstructures at the same four macroscopic states of interest as before. For interpreting the figures, we note that the diagonal line indicates perfect correlation (“1-to-1”)—in other words, the histogram bins which fall on this line indicate the counts of elements which have the same von Mises stress in both the gradient simulation and the homogenized simulation. Histogram bins above the line indicate counts of elements which had a higher von Mises stress in the gradient sample, while those below the line indicate the opposite. White regions in the plots are bins that have no counts of elements which exhibit the given stress pairing between the gradient and homogenized microstructures. We observe, qualitatively, that as plasticity begins and evolves, the distribution of element counts tends to skew toward a positive deviation from the line of perfect correlation, indicating that more elements exhibit higher von Mises stresses in the gradient virtual microstructure compared to the homogenized microstructure.